asset 1 has an expected return of 10% and a standard deviation of 20%. Asset 2 has an expected return of 15% and a standard deviation of 30%. the correlation between the two assets is -1.0. portfolios of these two assets will have a standard deviation of what?
Answer #1

Portfolio Variance = (wa1)^2 * SDa1 + (wa2)^2 * SDa2 + 2*(wa1)*(wa2)*Cor(a1, a2)

where w = weight, a = asset, SD = Standard deviation and Cor = Correlation between a1 & a2

Portfolio Variance = (0.5^2 * 0.2) + (0.5^2 * 0.3) + (2 * 0.5 *0.5 * -1.0) = 0.05 + 0.075 - 0.5 = 0.075

Standard Deviation of portfoliio = (0.075)^0.5 = 0.2739 or 27.39%